Chelsea Fashions is expected to pay an annual dividend of $1.26 a share next year. The market price of the stock is $24.09 and the growth rate is 2.6 percent. What is the cost of equity?

Multiple Choice

  • 7.91 percent

  • 9.77 percent

  • 9.24 percent

  • 7.54 percent

  • 7.83 percent
Answer #1

Here

Dividend next year D1 = $1.26

share price P= $24.09

growth rate g= 2.6%

cost of equity r= ?

we will use dividend discount model to find out cost of equity

share price P = D1/(r-g)

24.09 = 1.26/(r-2.6%)

24.09*r-24.09*0.026 =1.26

24.09*r = 1.26+0.626

24.09r = 1.89

r= 0.0783

r = 7.83%

Hence cost of equity is 7.83% so last option is correct here.

Thanks